The Higher Education Research Institute at UCLA collected data from 203,967 incoming first-time, full-time freshmen from 270 four-year colleges and universities in the U.S. 71.3% of those students replied that, yes, they believe that same-sex couples should have the right to legal marital status. Suppose that you randomly select freshman from the study until you find one who replies "yes." You are interested in the number of freshmen you must ask. What is the probability that you will need to ask fewer than three freshmen? (Round your answer to four decimal places.)
Answer #1

Ans:

Use Geometric distribution with p=0.713

P(x=k)=(1-0.713)k-1*0.713

P(x<3)=P(x=1)+P(x=2)

=0.713+(1-0.713)*0.713

= 0.9176